# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v2

@@ 5 Further results
To further understand the language model’s behaviour to prompts, we study whether the language model is respecting our in-context model class specifications (Appendix A.3) and test how the detail of the task description affects the posterior performance of the elicited priors (Appendix A.10) and how experts can supplement the process. Further to this, we explore how the number of task descriptions affects the priors elicited (Appendix A.11).
Finally, to contextualise the value of prior elicitation from language models, we report the cost of performing prior elicitation (Appendix A.12) and evaluate the quality of priors elicited from other language models with different price points (Appendix A.13.1) as well as open-source models (Appendix A.13.2).
@@ 6 Conclusion
In this work, we presented a method for extracting task information from an LLM to build probabilistic linear models with fewer data points. By asking a language model to play the role of a domain expert, we were able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples (Section 3.1). As also demonstrated in Appendix A.6.2, these prior distributions can significantly impact the delivery of care in healthcare settings, where collecting labels can be costly and time-consuming, but the need for accurate predictive models is great.
@@
Since our approach to prior elicitation can also be seen as a direct replacement for in-context learning for datasets with numerical features, we present a framework for deciding whether, given a task, to use prior elicitation or in-context learning based on the Bayes factor (Section 3.6). However, the Bayes factor assumes that the in-context learner is performing Bayesian inference. Therefore, we use maximum likelihood estimation and sampling distributions to approximate the internal model used during in-context learning and its prior and posterior distributions (Section 3.3). This allows us to inspect the truthfulness of the language model when we conduct prior elicitation (Section 3.4), as well as empirically question whether the language model is performing Bayesian inference on a given task (Section 3.5).
We find that prior elicitation is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error. Prior elicitation also extends the use of the language model’s prior knowledge of predictive tasks to previously infeasible situations, where feature observations might be frequent but labelled examples are few; where in-context learning is too slow or costly.
@@ A.1 Availability of code
A.2 Extracting internal prior and posterior
@@
Which corresponds to the maximum likelihood estimate (MLE) of $\varphi$ over some dataset $\hat{\mathcal{D}}$ given the task $T$. Now consider that $\hat{\mathcal{D}}$ is generated by a given $\varphi$ sampled from $\varphi\sim p_{M,T}(\varphi)$ and $X$. By sampling $x\sim\mathcal{U}[X]$ from a uniform distribution over $X$, we can create $S$ different datasets of the form:
@@ A.3 Does the language model use the model class we specify?
However, for all other prior extraction tasks presented in Section 3.3, we find that the language model uses a linear model in-context, demonstrated by the low MSE loss values between the predictions of our MLE samples and the true in-context predictions, shown in Figure 11. This is often not the case when the language model is presented with demonstration samples, as in the posterior extraction presented in Figure 5, in which the language model uses a linear combination of features to predict the targets for Heart Disease and Wine Quality only.
@@ A.4 Energy distance and statistic
Here ${M_{XY}}_{i,j}=||x_{i}-y_{j}||_{2}$ is the pairwise distance matrix.
@@ A.5.1 Dataset details
  • A private dataset collected as part of a study on Dementia care: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. In this dataset, the labels corresponded to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
@@
  • The California Housing dataset [37]: This dataset contains $20640$ examples of house values in California districts and $8$ features detailing information from the U.S census data for the property’s region. The goal of the dataset is to predict the median house for that region. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Wine Quality dataset [38, 39]: This dataset contains $6497$ data points of features relating to physicochemical characteristics of red and white wines with the goal of predicting whether the wines are of good quality or not. This dataset contained $2384$ examples of bad quality wine and $4113$ examples of good quality wine. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
Where datasets are publicly available, we perform minimal processing for ease of reproducibility. In some cases, we replace the feature names with ones that are more descriptive (taken from the dataset descriptions also available publicly) to aid the language model’s comprehension of the data. All of these changes to the raw data are available within the supplementary code, which will automatically download and process the data upon loading.
@@ A.5.2 Evaluating prior elicitation
To evaluate the usefulness of the priors elicited from the language model (presented in Section 2.2), we first studied their posterior performance in comparison to using an uninformative prior. To do this, we chose the tasks outlined in Section 2.5, $y=2x_{1}-x_{2}+x_{3}$, Breast Cancer, California Housing, Urinary Tract Infection (UTI), Wine Quality, and Heart Disease and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
For the UTI task, we split the dataset so that each participant’s data is only in either the testing or training set, as some participants produced multiple examples. For the other classification tasks, the splits are stratified using the targets. To test the posterior performance after observing the given number of data points in the experiments, we randomly select examples from the training set.
After performing prior elicitation (Section 3.1), for each of the tasks, we have $100$ Gaussian priors for each feature in the dataset. We then build a mixture of the priors using Equation 1, where we weigh the Gaussian distribution in each dimension and from each call to the language model separately.
@@
To produce the results in Figures 2(a) and 8, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. These metric values are then presented in Figures 2(a) and 8, which show their histograms. To produce Figures 2(b) and 13, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting their mean and standard deviation.
For the synthetic task, one of the example task descriptions is as follows:
@@ A.5.3 Extracting the in-context model’s hidden prior and posterior
For each of these linear models, we then calculate the regression labels or the classification logits using the same data given to the language model for in-context predictions. This allows us to calculate an error in our estimation of the in-context model, which depends on how well the in-context learner is approximating a linear model.
@@
We can then measure, as in Section 3.4, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.4) between this sample and the extracted in-context sample. This is done to test whether the language model is using the same distribution we elicit from it as the one it uses in-context. The results of this are presented in Table 1.
@@
To do this, we start by fitting a Gaussian kernel density estimator (with a bandwidth factor of $0.25$) to our in-context model prior parameter samples. We then use the No-U-Turn sampler [35] with $1000$ adaption steps and the same demonstration examples given during in-context learning to sample posterior parameters. Here, we use $100$ chains with $10000$ samples each to build an approximate posterior distribution.
We can then use the energy statistic to measure the difference between the extracted in-context posterior distribution and the Monte Carlo approximation of the posterior based on the extracted in-context prior distribution.
The values presented in Table 2 show the result of this experiment.
@@ A.5.4 Calculating Bayes factor
For a single dataset, we start by randomly selecting $25$ data points, $5$ times, for which we consider the prior predictive distribution. For prior elicitation this involves sampling the prior predictive distribution (drawing $500$ samples for each split), and for in-context learning this involves prompting a language model for a prediction for each of the $100$ task descriptions (producing $100$ samples for each of the splits).
For both sets of prior predictive samples, we calculate their log-likelihood using either cross-entropy loss or mean squared error, depending on whether the task is classification or regression. For each dataset, for each test split, and for each method, we calculate the mean of this log-likelihood over the samples. With these values, for each dataset and test split, we measure the difference between the value for the elicited prior predictions and the in-context predictions to get the log Bayes factor.
We then report the mean Bayes factor over the splits along with the standard deviation in Figure 6, which helps us to decide, given a dataset, which method is the most appropriate.
@@ A.5.5 Given a new dataset
    Ask a language model to rephrase your user role and system role many times. In our experiments, we asked the language model to rephrase both descriptions $10$ times for each.
@@
    Use these individual Gaussian priors to build a mixture of Gaussians as a prior for a linear model.
@@
Now, to measure whether prior elicitation or in-context learning is better for a given task (assuming the language model is approximating Bayesian inference well):
@@
    Then, to perform in-context learning, use a similar task description to above but with a different system role, that specifies the language model should make predictions instead of providing prior distributions. With this new set of descriptions, again ask a language model to rephrase them.
@@
    To sample the prior predictive, using the product of these descriptions, ask a language model to make predictions on the data directly.
@@
    Measure the Bayes factor between the likelihood of the prior predictive distribution given using prior elicitation and the in-context model on the already collected data.
@@
    Using the Bayes factor, decide whether in-context learning or prior elicitation is more appropriate.
To empirically check whether the language model is approximating Bayesian inference in its in-context predictions or to test whether it is being faithful to the priors we elicit, we can follow the below procedure:
@@
Therefore, for the latter, we can follow the below procedure:
@@
    Use a method to calculate the difference between two samples of a distribution to understand the difference between the posterior samples calculated through Monte Carlo sampling of the in-context prior and the samples of the in-context posterior.
This last procedure might indicate that our in-context learner is not approximating Bayesian inference and is not suitable for a given task, where it could be important to us that our predictive model is updating its prior knowledge in an intuitive way.
@@ A.6 Further prior elicitation results
Figure 8: Test accuracy of the posterior distribution over the model weights of a logistic and linear regression model for varied numbers of training data points. We evaluate the accuracy and mean squared error of different prior distributions over the weights of a logistic and linear regression model. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $100$ sampled priors and $5$ random splits of the training and testing data. Note that the bottom row shows the mean squared error (on log-scale) where lower values are better. The green arrows point in the direction of improvement in the metric.
Here, we present the results of the same experiment as discussed in Section 3.1, applied to two other datasets discussed in Section 2.5.
The results of a logistic regression model on a test set after observing training data and using a prior elicited from the language model or using an uninformative prior are shown in Figure 8.
These results suggest that for the Wine Quality dataset, the prior elicited is better than the uninformative prior for smaller numbers of observations but is too strict when the number of observations is larger since the result distribution is more rigid than for an uninformative prior. This is similar to what is observed in the results on the Breast Cancer and UTI datasets presented in Section 3.1.
Furthermore, for the Heart Disease dataset, we find that the prior elicited from the language model performed comparably to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the language model was unsure about the prior distribution on the parameters for a model predicting heart disease. The parameter distributions elicited can be seen in Figure 11, which contextualises these results.
@@ A.6.2 UTI accuracy with the label collection context
It is important to note that the accuracy in Figure 10 is not achieved as a result of training on the earliest collected data points. This is because in our analysis we perform cross-validation and use a subset of the data from a later time period (Section A.5.1). There were changes to the in-home devices during the study, such as the addition of a device to collect sleep information, which we use within the dataset evaluated in the main results. However, if all data collection devices had been used from the beginning of the study, we could expect the accuracy in time to resemble the figure presented here.
A.7 Is the language model being truthful when providing priors through elicitation?
@@
Figure 11 visualises the different prior parameter distributions for the UTI, Breast Cancer, California Housing, Wine Quality, and Heart Disease datasets in each feature dimension. Notable are the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 3.4. In the in-context case, prior distributions frequently have modes close to 00, suggesting that the internal model is not sure about the relationship between features and targets. However, the elicited priors often represent a stronger belief in the correlation between features and targets, with most histograms achieving modes greater than or less than 00.
@@
Furthermore, the prior distribution on the parameters for the Heart Disease dataset have more variation in their values in comparison to that extracted from the language model’s internal model. This tells us that the elicited prior parameters represent more uncertainty in their value and are therefore more flexible when calculating the posterior on observed data.
A.8 Testing empirically whether the language model is approximating Bayesian inference
In this section, we visually inspect the differences between the posterior distribution extracted through maximum likelihood sampling of the language model’s internal model (Section 3.3) and one that is sampled through MCMC methods on the extracted prior distribution.
Figure 12: Distribution of the posterior parameters extracted from the internal model for the predictive tasks against the posterior MC samples with extracted prior for a single data split. Here, we compare the distribution of the posterior extracted from the language model against the distribution that we sample using MCMC methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined. This is calculated using one of the $5$ training sets tested, hence the difference in values compared to Table 2.
For a single random draw of $25$ training data points from each corresponding dataset, in Figure 12, we sample parameter values from the posterior distributions and present them as a histogram in each feature dimension.
In the cases of Breast Cancer, Wine Quality, and Heart Disease, we see that the posterior distributions overlap significantly and show a strong agreement, which is confounded by the small value of the energy statistic between the histograms calculated across all features combined. This is empirical evidence that the language model is performing Bayesian inference in-context, since it is aligned with the posterior estimates from an MCMC method.
On the California Housing dataset, the posteriors do not match as closely. The histograms given in Figure 12, together with the comparatively larger energy distance, show a significant difference in distributions. In the extracted posterior, almost all of the values are contained within a small range around zero, whilst the posterior sampled from MCMC methods shows a larger variation in possible values. Additionally, in these cases, the energy statistic is far greater than for other datasets and suggests that, for the California Housing regression task, in-context learning is not approximating Bayesian inference.

### v3

@@ 5 Further results
To further understand the language model’s behaviour to prompts, we study whether the language model is respecting our in-context model class specifications (Appendix A.3) and test how the detail of the task description affects the posterior performance of the elicited priors (Appendix A.10), as well as how experts can supplement the process in a controlled synthetic experiment. Further to this, we explore how the number of task descriptions affects the priors elicited (Appendix A.11).
Finally, to contextualise the value of prior elicitation from language models, we report the cost of performing prior elicitation (Appendix A.12) and evaluate the quality of priors elicited from other language models with different pricing tiers (Appendix A.13.1) as well as open-source models (Appendix A.13.2).
@@ 6 Conclusion
In this work, we presented a method for extracting task information from an LLM to build probabilistic linear models with fewer data points. By asking a language model to play the role of a domain expert, we were able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples (Section 3.1). As demonstrated in Appendix A.6.2, these prior distributions can significantly impact the delivery of care in healthcare settings, where label collection can be costly and time-consuming, but the need for accurate predictive models is great.
@@
Since our approach to prior elicitation can also be seen as a direct replacement for in-context learning for datasets with numerical features, we present a model selection framework, based on the Bayes factor (Section 3.6). However, the Bayes factor assumes that the in-context learner is performing Bayesian inference. Therefore, we use maximum likelihood estimation and sampling distributions to approximate the internal model used during in-context learning and its prior and posterior distributions (Section 3.3). This allows us to inspect the truthfulness of the language model when we conduct prior elicitation (Section 3.4), as well as empirically question whether the language model is performing Bayesian inference on a given task (Section 3.5).
We find that prior elicitation is an improved alternative to in-context learning in our setting, as we can inspect priors, confidently perform Bayesian inference, and often achieve reduced error. Prior elicitation also extends the use of the language model’s prior knowledge of predictive tasks to previously infeasible situations, such as low-resource environments or where feature observations are frequent but labelled examples are few.
@@ A.1 Availability of code
A.2 Formal details on extracting the internal prior and posterior
@@
This corresponds to the maximum likelihood estimate (MLE) of $\varphi$ over some dataset $\hat{\mathcal{D}}$ given the task $T$. Now consider that $\hat{\mathcal{D}}$ is generated by a given $\varphi$ sampled from $\varphi\sim p_{M,T}(\varphi)$ and $X$. By sampling $x\sim\mathcal{U}[X]$ from a uniform distribution over $X$, we can create $S$ different datasets of the form:
@@ A.3 Does the language model use the model class we specify?
However, for all other prior extraction tasks presented in Section 3.3, we find that the language model uses a linear model in-context, demonstrated by the low MSE loss values between the predictions of our MLE samples and the true in-context predictions, shown in Figure 12. This is often not the case when the language model is presented with demonstration samples, as in the posterior extraction presented in Figure 5, in which the language model uses a linear combination of features to predict the targets for the Heart Disease and Diabetes tasks only.
@@ A.4 Energy distance and statistic
Here, ${M_{XY}}_{i,j}=||x_{i}-y_{j}||_{2}$ is the pairwise distance matrix.
@@ A.5.1 Dataset details
  • A private dataset collected as part of a study on Dementia care  [^3] https://www.imperial.ac.uk/uk-dri-care-research-technology/: This dataset contains $10$ features engineered from measurements of in-home activity and physiology. In this dataset, the labels corresponded to days with a clinical labelling of positive or negative urinary tract infection (UTI) which were calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of the data containing $78$ people and $371$ labelled days of UTI, with $270$ negatives and $101$ positives. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Hypothyroid dataset [39, 40]: This dataset, processed from the original, consists of $150$ examples of patients’ thyroid-related blood levels and whether they have a hypothyroidism diagnosis or not. This dataset contains $4$ features, with $75$ examples of a patient with hypothyroidism and $75$ examples of a patient without hypothyroidism. When training and testing on this dataset, we centred and scaled the training data and used the same statistics to normalise the test data.
  • The Diabetes dataset [38]: After preprocessing, this dataset contains $442$ people with Diabetes and some physiological measurements. Along with some demographic information, the task is to predict a quantitative measure of their diabetes disease progression one year after this baseline. This dataset consisted of $10$ features and the targets were regression values with a maximum value of $10.0$, minimum value of $0.0$, and a mean value of $3.96$.
@@
Where datasets are publicly available, we perform minimal processing for ease of reproducibility. In some cases, we replace the feature names with ones that are more descriptive (taken from the dataset descriptions also available publicly) to aid the language model’s comprehension of the data. All of these changes to the raw data are available within the supplementary code, which will automatically download and process the data when loading.
@@ A.5.2 Evaluating prior elicitation
To evaluate the usefulness of the priors elicited from the language model (presented in Section 2.2), we started by comparing their posterior performance to an uninformative prior. To do this, we chose the predictive tasks described in Section 2.5: $y=2x_{1}-x_{2}+x_{3}$, Urinary Tract Infection (UTI), Heart Disease, Hypothyroidism, Diabetes, and Breast Cancer and calculated the posterior accuracy or mean squared error on a test set after observing varied numbers of examples. For each of the datasets, we selected $50\%$ of data points to form the test set and used the remaining data points to form a training set. This is done $10$ times to produce $10$ folds of random training and testing splits.
For the UTI task, we split the dataset so that each participant’s data is only in either the testing or training set, as some participants produced multiple examples. For the other classification tasks, the splits are stratified using the targets. To evaluate the posterior performance after observing a given number of data points, we randomly select that number of examples from the training set to train on and then test on the same test set as before.
After performing prior elicitation (Section 3.1), for each of the tasks, we have $100$ Gaussian priors for each feature in the dataset. We then build a mixture of the priors using Equation 1, where we weight the Gaussian distribution in each dimension and each task description separately.
@@
To produce the results in Figures 2(a) and 9, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. To produce Figures 2(b) and 14, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting their mean and standard deviation.
As an example, for the synthetic task one of the task descriptions is as follows:
@@ A.5.3 Extracting the in-context model’s hidden prior and posterior
For each of these linear model samples, we calculate the regression labels or the classification logits using the same data given to the language model for in-context predictions. This allows us to calculate an error in our estimation of the in-context model, which depends on how well the in-context learner is approximating a linear model.
@@
We can then measure, as in Section 3.4, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.4) between this sample and the extracted in-context sample and through visual inspection. By doing this, we test whether the language model is using the same distribution we elicit from it as the one it uses in-context. The results of this are presented in Table 1.
@@
To do this, we start by fitting a Gaussian kernel density estimator (with a bandwidth factor of $0.25$) to our in-context model prior parameter samples. We then use the No-U-Turn sampler [35] with $1000$ adaption steps and the same demonstration examples given during in-context learning to sample posterior parameters. Here, we use $100$ chains of $10000$ samples to build an approximate posterior distribution.
We can then use the energy statistic to measure the difference between the extracted in-context posterior distribution and the Monte Carlo approximation of the posterior based on the extracted in-context prior distribution. The values presented in Table 2 show the result of this experiment.
@@ A.5.4 Calculating Bayes factor
For a single dataset, we start by randomly selecting $25$ data points, $5$ times, and consider the prior predictive distribution. For prior elicitation this involves sampling the prior predictive distribution (drawing $500$ samples for each split), and for in-context learning this involves prompting a language model for a prediction for each of the $100$ task descriptions (producing $100$ samples for each of the splits).
For both sets of prior predictive samples, we calculate their log-likelihood using either cross-entropy loss or mean squared error, depending on whether the task is classification or regression. For each dataset, for each test split, and for each method, we calculate the mean of this log-likelihood over the samples. With these values, we measure the difference between the value for the elicited prior predictions and the in-context predictions to obtain the log Bayes factor.
We then report the mean and standard deviation of the Bayes factor over the splits in Figure 7, which helps us to decide, given a dataset, which method is the most appropriate; assuming that in-context learning approximates Bayesian inference.
@@ A.5.5 Given a new dataset
    Ask a language model to rephrase the user role and system role many times. In our experiments, we asked the language model to rephrase both descriptions $10$ times for each.
@@
    Use these individual Gaussian priors to build a mixture of Gaussians as a prior distribution for a linear model.
@@
Functions for doing this are provided in the supplementary code.
Now, to measure whether prior elicitation or in-context learning is better for a given task (assuming the language model is approximating Bayesian inference):
@@
    Then, to perform in-context learning, use a similar task description to above but with a different system and user role that specifies the language model should make predictions instead of providing prior distributions. With this new set of descriptions, again ask a language model to rephrase them.
@@
    To sample the prior predictive distribution using the product of these descriptions, ask a language model to make predictions on the data directly.
@@
    Measure the Bayes factor between the likelihood of the prior predictive distribution using prior elicitation and the in-context model on the already collected data.
@@
    Using the Bayes factor and by factoring in resource cost, decide whether in-context learning or prior elicitation is more appropriate.
To empirically check whether the language model is approximating Bayesian inference to produce its in-context predictions or to test whether it is being faithful to the priors we elicit, we can follow the below procedure:
@@
For this, we can follow the below procedure:
@@
    Use a method to calculate the difference between two samples of a distribution to understand the differences between the posterior samples calculated through Monte Carlo sampling of the in-context prior and the samples of the in-context posterior.
This last procedure might indicate that our in-context learner is not approximating Bayesian inference and is not suitable for a given task, since it could be important to us that our predictive model is updating its prior knowledge in a reliable way.
@@ A.6 Further prior elicitation results
Figure 9: Test accuracy of the posterior distribution over the model weights of a logistic and linear regression model for varied numbers of training data points. We evaluate the accuracy and mean squared error of different prior distributions over the weights of a logistic and linear regression model. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. The green arrows point in the direction of improvement in the metric.
In Figure 9 we present the results of the same experiment as discussed in Section 3.1, applied to three other datasets discussed in Section 2.5.
For the synthetic task $y=2x_{1}-x_{2}+x_{3}$, the priors elicited by LLM allow for a posterior predictive distribution with a markedly reduced mean squared error over an uninformative prior and often produce less variation in the result. Due to the simple nature of the task however, after only a few training examples, both priors achieve similar posterior results.
In the case of Hypothyroid prediction, the LLM-elicited prior produces similarly accurate posterior parameter samples than the uninformative prior for smaller numbers of observations. However, as the number of observations increases, the LLM-elicited prior begins to produce more accurate posterior distributions.
For the Heart Disease dataset, we find that the prior elicited from the language model performed similarly to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the language model was unsure about the prior distribution on the parameters for a model predicting heart disease. The parameter distributions elicited can be seen in Figure 12, which contextualises these results.
@@ A.6.2 UTI accuracy with the label collection context
It is important to note that the accuracy in Figure 11 is not achieved as a result of training on the earliest collected data points. This is because in our analysis we perform cross-validation and use a subset of the data from a later time period (Section A.5.1). Changes in the in-home devices were made during the study, such as the addition of a device to collect sleep information, which we use within the dataset evaluated in the main results. However, if all data collection devices had been used from the beginning of the study, we could expect the accuracy in time to resemble the figure presented here.
A.7 Is the language model being truthful when we elicit priors?
@@
Figure 12 visualises the different prior parameter distributions for all of the tested datasets in each feature dimension. Notable are the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 3.4. In the in-context case, prior distributions often have modes close to 00, suggesting that the internal model is not sure about the relationship between features and targets. The elicited priors often represent a stronger belief in the correlation between features and targets, with most histograms achieving modes greater than or less than 00.
@@
Furthermore, the prior distributions on the parameters for the Heart Disease and Hypothyroid datasets have more variation in their values in comparison to that extracted from the language model’s internal model. This tells us that the elicited prior parameters represent more uncertainty in their value and are therefore more flexible when calculating the posterior on observed data.
A.8 Empirically, is the language model approximating Bayesian inference?
In this section, we present the differences between the posterior distribution extracted through maximum likelihood sampling of the LLM’s in-context internal model (Section 3.3) and one that is sampled through MC methods on the extracted in-context prior distribution.
Heart Disease Hypothyroid Diabetes Breast Cancer
0.04 ± 0.01 0.06 ± 0.01 0.02 ± 0.00 0.01 ± 0.00
Figure 13: The different posterior samples Here, we compare the distribution of the posterior extracted from the language model’s in-context predictions against the distribution that we sample using Monte Carlo methods on the extracted prior distribution. Both posterior distributions are based on the same $25$ data point sample from the corresponding dataset. We draw the histogram separately in each dimension, but calculate the energy distance across all features combined.
Table 2 shows the mean and standard deviation of the energy distance between these distributions, for all of the datasets tested, including those that have a high approximation error when we estimate the in-context posterior. In addition, Figure 13 shows the distribution of these posterior parameters for visual inspection.
On the other hand, for the Hypothyroid predictive task, the parameter distributions differed significantly, with the in-context prior having a less varied distribution. This is reflected in the energy distance, which is the largest of all of the datasets tested.
These comparisons must be made with the approximation error in mind, presented in Section 3.3, which suggests that the posterior in-context distributions for the Breast Cancer and Hypothyroid tasks are not reliable.